OpenStudy (anonymous):

I need help with a question from a project I'm doing. I finished the rest of the assignment but the last question makes no sense to me. It's a simulation lab online but here is the last question.

OpenStudy (anonymous):

Place the rider at Point E and start his motion. Point E is 24 meters high. The mass of the rider was given in question 4 as 100 kg. Assume that point B touches the ground. Using the equations and the principle of conservation of energy, determine the rider’s kinetic energy at point B. (Hint: You do not know his velocity at the point, there is another way to figure it out)

OpenStudy (shamim):

Given m=100 kg h=24m g=9.8m/s^2

OpenStudy (shamim):

Potential energy at point E P.E=mgh=?

OpenStudy (shamim):

U know potential energy at point E= kinetic energy at point B=?

OpenStudy (shamim):

Response plz for getting better help!!!

OpenStudy (anonymous):

How do you get 9.8?

OpenStudy (anonymous):

m=100 kg h=24m g=9.8m/s^2 PE = mgh PE = (100kg)(9.8m/s^2)(24m) PE = 23520 J K = 1/2(mv)^2 K = 1/2(24kg)(9.8m/s)^2 K = 1,152.48 J

OpenStudy (shamim):

g=acceleration due to gravity

OpenStudy (shamim):

The value of g is a constant

OpenStudy (shamim):

g=9.8m/s^2

OpenStudy (shamim):

So g is not a velocity

OpenStudy (anonymous):

So how do I find out velocity from the info given?

OpenStudy (shamim):

Potential energy at point E=mgh=100*9.8*24=?

OpenStudy (shamim):

Kinetic energy at point B=100*9.8*24=?

OpenStudy (shamim):

So .5mv^2=100*9.8*24 .5*100*v^2=100*9.8*24 v=?